Through a refinery, fuel ethanol is flowing in a pipe at a velocity of 1 m/s and a pressure of 101300 Pa. The refinery needs the ethanol to be at a pressure of 2 atm (202600 Pa) on a lower level. The density of ethanol is 789 kg/m3 and gravity g is 9.8 m/s2. How far must the pipe drop in height in order to achieve this pressure? Assume the velocity does not change.


1
Expert's answer
2021-02-24T12:46:00-0500

From the Bernoulli's equation:


"p_1+\frac{\rho v_1^2}{2}+\rho gh_1=p_2+\frac{\rho v_2^2}{2}+\rho gh_2,"


where "p_1" and "p_2" - the pressures at the levels h1 and h2 respectively;

"v_1=v_2=v" - the velocity at the same levels;

"\rho" - the density of ethanol;

g - the gravitational acceleration.


So,


"h_2-h_1=\frac{p_1-p_2}{\rho g}=\frac{202600-101300}{789\cdot 9.8}=13.1\space m."


Answer: 13.1 m.